OpenStudy (anonymous):

bit harder question. The ratio of wheat and rice in conatiner A is 2 : 3 and that in box B is 3 : 7. What quantities must be taken from A and B (in kg resp) to form a mixture containing 5 kg of wheat and 11 kg of rice?

OpenStudy (dumbcow):

wheat: 2A +3B = 5 rice : 3A +7B = 11 solve system for A,B using elimination method

OpenStudy (anonymous):

sorry b=7/5

OpenStudy (anonymous):

a=2/5,b=7/5

OpenStudy (dumbcow):

wait sorry i was wrong in setting it up .... your answers are correct though :) however it doesn't make sense

OpenStudy (dumbcow):

(2/5)A + (3/10)B = 5 (3/5)A + (7/10)B = 11 A + B = 16

OpenStudy (anonymous):

a=2 b=14

OpenStudy (anonymous):

i have 5 to 10 questions can u help me

OpenStudy (dumbcow):

ok